# On the proximal dynamics between integrable and non-integrable members of a generalized Korteweg-de Vries family of equations

### v1

@@ Abstract.
The distance between the solutions to the integrable Korteweg-de Vries (KdV) equation and a broad class of non-integrable generalized KdV (gKdV) equations is estimated in appropriate Sobolev spaces. Special cases of the latter family of equations include the standard gKdV equation with a power nonlinearity, as well as weakly nonlinear perturbations of the KdV equation. The distance estimates are established for initial data and nonlinearity parameters of arbitrary size. A crucial step for their derivation is the proof of a size estimate for local solutions to the gKdV family of equations, which is linear in the norm of the initial data. As a result, the distance estimates themselves predict that the dynamics between the gKdV and KdV equations remains close over long times for initial amplitudes even close to unity, while for larger amplitudes they predict an explicit rate of deviation of the dynamics between the compared equations. These theoretical results are illustrated via numerical simulations in the case of one-soliton and two-soliton initial data, which are in an excellent agreement with the theoretical predictions. Importantly, in the case of a power nonlinearity and large solitonic initial data, the deviation between the integrable and non-integrable dynamics can be drastically reduced upon incorporating suitable rotation effects by means of a rescaled KdV equation. As a result, the integrable dynamics stemming from the rescaled KdV equation may persist in the gKdV family of equations over remarkably long timescales.
@@ 2. Proof of the size estimate
In the case of $U_{0}\in\mathcal{S}(\mathbb{R})$, the results of [Kat83] (see also [Tem69, Sjö70, BS75]) imply the existence of a unique solution in $U\in C_{t}^{\infty}([0,T_{0}];\mathcal{S}(\mathbb{R}))$, where $T_{0}>0$ depends only on the norm of $U_{0}$. For that smooth solution, let $n\in\mathbb{N}$ with $0\leq n\leq s$ and apply $\partial_{x}^{n}$ to equation (1.1). Then, multiply the resulting equation by $\partial_{x}^{n}U$ and integrate in $x$ while recalling that $U$ is real-valued to obtain
@@
For $0\leq j\leq n-2$, by the Sobolev embedding theorem
$\left\|\partial_{x}^{n-j}(F(U))\,\partial_{x}^{j+1}U\right\|_{L_{x}^{2}(\mathbb{R})}\leq\left\|\partial_{x}^{j+1}U\right\|_{L_{x}^{\infty}(\mathbb{R})}\left\|\partial_{x}^{n-j}(F(U))\right\|_{L_{x}^{2}{\mathbb{R}}}\leq c\left\|U\right\|_{H_{x}^{j+2}(\mathbb{R})}\left\|\partial_{x}^{n-j}(F(U))\right\|_{L_{x}^{2}{\mathbb{R}}}$ (2.12)
@@
$\left\|\partial_{x}^{n-j}(F(U))\,\partial_{x}^{j+1}U\right\|_{L_{x}^{2}(\mathbb{R})}\leq c\left\|U\right\|_{H_{x}^{s}(\mathbb{R})}\left\|\partial_{x}^{n-j}(F(U))\right\|_{L_{x}^{2}{\mathbb{R}}}.$ (2.13)
@@
Next, we extend the result to the case of Sobolev solutions. To this end, by density, given $U_{0}\in H^{s}(\mathbb{R})$, we let $\{U_{0,n}\}_{n\in\mathbb{N}}\in\mathcal{S}(\mathbb{R})$ be such that $\left\|U_{0}-U_{0,n}\right\|_{H_{x}^{s}(\mathbb{R})}\to 0$ as $n\to\infty$. Then, thanks to the result of [Kat83], there is a unique solution $U_{n}\in C_{t}^{\infty}([0,T_{0,n}];\mathcal{S}(\mathbb{R}))$ corresponding to $U_{0,n}$, where $T_{0,n}>0$ depends only on the norm of $U_{0,n}$. Hence, by estimate (1.21), which was established above for smooth solutions,
@@
$\leq 2\left\|U_{0,n}\right\|_{H^{2}(\mathbb{R})}+K\left\|U_{0}-U_{0,n}\right\|_{H^{s}(\mathbb{R})},\quad t\in[0,T_{n}].$
@@ 3. Proof of the distance estimates
$\left\|\Delta(t)\right\|_{L_{x}^{2}(\mathbb{R})}\leq\left\|\Delta(0)\right\|_{L_{x}^{2}(\mathbb{R})}+\int_{0}^{t}\left\|N(\tau)\right\|_{L_{x}^{2}(\mathbb{R})}d\tau.$ (3.6)
@@
$\left\|\Delta(t)\right\|_{L_{x}^{2}(\mathbb{R})}\leq\left\|\Delta(0)\right\|_{L_{x}^{2}(\mathbb{R})}+\sup_{\tau\in[0,t]}\Big[\left\|u(\tau)\right\|_{H_{x}^{1}(\mathbb{R})}^{2}+c_{k}\Big(\left\|U(\tau)\right\|_{H_{x}^{1}(\mathbb{R})}^{k+1}+B_{k}\sum_{j=1}^{k-1}\left\|U(\tau)\right\|_{H_{x}^{1}(\mathbb{R})}^{j+1}\Big)\Big]\,t.$ (3.10)
@@
$\sup_{t\in[0,T]}\left\|\Delta(t)\right\|_{L_{x}^{2}(\mathbb{R})}\leq\left\|\Delta(0)\right\|_{L_{x}^{2}(\mathbb{R})}+2^{k+1}\Big[\left\|u_{0}\right\|^{2}_{H^{2}(\mathbb{R})}+c_{k}\Big(\left\|U_{0}\right\|_{H^{2}(\mathbb{R})}^{k+1}+B_{k}\sum_{j=1}^{k-1}\left\|U_{0}\right\|_{H^{2}(\mathbb{R})}^{j+1}\Big)\Big]\,T.$ (3.11)
@@
$\left\|\Delta(t)\right\|_{H_{x}^{s}(\mathbb{R})}\leq\sqrt{2}\left\|\Delta(0)\right\|_{H_{x}^{s}(\mathbb{R})}+\sqrt{2}\int_{0}^{t}\left\|N(\tau)\right\|_{H_{x}^{s}(\mathbb{R})}d\tau.$ (3.13)
@@
$\left\|\Delta(t)\right\|_{H_{x}^{s}(\mathbb{R})}\leq\sqrt{2}\left\|\Delta(0)\right\|_{H_{x}^{s}(\mathbb{R})}+\sqrt{2}\sup_{\tau\in[0,t]}c_{s,k}\Big(\left\|u(\tau)\right\|_{H_{x}^{s+1}(\mathbb{R})}^{2}+\left\|U(\tau)\right\|_{H_{x}^{s+1}(\mathbb{R})}^{k+1}+B_{k}\sum_{j=1}^{k-1}\left\|U(\tau)\right\|_{H_{x}^{s+1}(\mathbb{R})}^{j+1}\Big)\,t$ (3.17)
@@
$\sup_{t\in[0,T]}\left\|\Delta(t)\right\|_{H_{x}^{s}(\mathbb{R})}$ $\leq\sqrt{2}\left\|\Delta(0)\right\|_{H_{x}^{s}(\mathbb{R})}+2^{k+\frac{3}{2}}\Big(\left\|u_{0}\right\|_{H^{s+1}(\mathbb{R})}^{2}+\left\|U_{0}\right\|_{H^{s+1}(\mathbb{R})}^{k+1}+B_{k}\sum_{j=1}^{k-1}\left\|U_{0}\right\|_{H^{s+1}(\mathbb{R})}^{j+1}\Big)\,T.$ (3.18)
@@ 4.1. The gKdV with power nonlinearity
As a first example, we consider the initial datum $u_{\text{s}}(x)$ in (4.4) with $c=0.01$ so that, according to (4.5), $\left\|u_{\text{s}}\right\|_{H^{2}}\approx 0.1551$. We then evolve this datum under the supercritical gKdV equation (1.3) with $k=5$ over the time interval $t\in[0,\left\|u_{\text{s}}\right\|_{H^{2}}^{-1}]\simeq[0,6.5]$ and hence compute numerically the difference norm $\left\|\Delta(t)\right\|_{\mathcal{X}}$ for $\mathcal{X}=H_{x}^{s}(\mathbb{R})$, $s=0,1$, and $L_{x}^{\infty}(\mathbb{R})$. By Corollary 1, we have the theoretical bound $\left\|\Delta(t)\right\|_{\mathcal{X}}\lesssim\epsilon^{2}\approx 0.024$. The numerical simulations fulfill this proximity estimate (see Figure 1) and, in fact, for $t\in[0,80]$ yield $\left\|\Delta(t)\right\|_{L_{x}^{\infty}}\leq 0.0024$, which is an order of magnitude smaller than the theoretical bound. Moreover, the simulations exhibit a linear in time growth of the proximity error, in agreement with the theoretical predictions.
Figure 1. The solution $U(x,t)$ of the gKdV equation (1.1) with monomial nonlinearity according to (4.1). Top left: The supercritical case $k=5$ supplemented with KdV one-soliton initial data $u_{\text{s}}(x)$ with $c=0.01$ and $x_{0}=0$. Top right: The critical case $k=4$ supplemented with mKdV one-soliton initial data $U_{\text{s}}(x)$ with $c=0.004$ and $x_{0}=0$. Bottom: The corresponding time evolutions of the norms $\left\|\Delta(t)\right\|_{\mathcal{X}}$ of the difference between the above gKdV solutions and the corresponding KdV and mKdV one-solitons for $\mathcal{X}=H_{x}^{1}$, $L_{x}^{2}$ and $L_{x}^{\infty}$.
As a second example, we consider the second initial profile $U_{\text{s}}(x)$ in (4.4) with $c=0.004$ so that, by (4.5), $\left\|U_{\text{s}}\right\|_{H^{2}}\approx 0.8718$. We evolve the mKdV one-soliton data under the critical gKdV equation (1.3) with $k=4$. For $t\in[0,\left\|U_{\text{s}}\right\|_{H^{2}}^{-2}]\approx[0,1.32]$, we compute numerically $\left\|\Delta(t)\right\|_{\mathcal{X}}$ (see Figure 1). By Corollary 1, we have $\left\|\Delta(t)\right\|_{\mathcal{X}}\lesssim\epsilon^{2}\approx 0.76$. Once again, the numerical results fulfill the theoretical estimates. Specifically, we find $\left\|\Delta(t)\right\|_{L_{x}^{\infty}}\leq 0.01$ for $t\in[0,100]$, which is again about an order of magnitude smaller than the theoretical bound. As in the first example, the growth of the distances in Figure 1 is in agreement with the theoretically predicted linear temporal growth of the proximity estimates.
@@
Figure 2. Left and middle: Snapshots comparing the profile of the solution of gKdV with power nonlinearity (1.3) for $k=3$ (solid blue) against the profile of the solution of the integrable KdV (1.4) with (red dashed). Right: The top panel shows the contour plot of the evolution of the initial datum $u_{\text{s}}(x)$ in (4.4) with $c=0.4$ and $x_{0}=0$ under gKdV. The dashed black line is the caustic of the analytical KdV soliton. The bottom panel presents the evolution of the distance norms between the solutions $\mathcal{X}=H_{x}^{s}$, $s=0,1,2$, and $L_{x}^{\infty}$.
The first two columns in Figure 2 are snapshots at specific times comparing the profiles of the solutions. The gKdV solution is shown with the solid blue curve and the KdV soliton with the dashed red curve. In the third column, the top panel shows the contour plot of the gKdV solution, together with the caustic of the corresponding analytical KdV soliton, depicted by the dashed black line. The bottom panel shows the evolution of the distance norm $\left\|\Delta(t)\right\|_{\mathcal{X}}$ with $\mathcal{X}=H_{x}^{s}(\mathbb{R})$, $s=0,1,2$, and $L_{x}^{\infty}(\mathbb{R})$.
According to Corollary 1, we expect proximity of the solutions only for very small timescales predicted by $\widetilde{T}_{s}=\mathcal{O}(\left\|u_{\text{s}}\right\|_{H^{2}}^{-3})$. This is already seen in the first snapshot at $t=1$, where we observe the beginning of the deviation of the numerical gKdV soliton from the analytical KdV one-soliton, manifested by the small increase in its amplitude and the onset of its radiation. In the snapshot at $t=10$, the radiation of the gKdV soliton is clearly observable and the KdV soliton is faster. These snapshots, as well as the one at $t=25$, depict profiles in the regime of the linear growth of the distances between the solutions. In particular, the snapshot at $t=25$ refers to a time close to the end of the linear growth regime of $\left\|\Delta(t)\right\|_{\mathcal{X}}$, manifested by the complete separation of the soliton profiles, as observed in the snapshot at $t=200$; the beginning of this separation at $t\sim 25$ is also visible in the contour plots of the solutions.
@@
We detect the following dynamical feature: for a given nonlinearity exponent $k\geq 2$, there exists a scaling parameter $0<\nu(k)<1$ such that the soliton dynamics in the gKdV equation remain remarkably proximal, for significantly large time intervals, to the analytical rescaled soliton (4.7).   For gKdV with $k=3$, we detect a scaling parameter $\nu(3)\sim 0.47$. Figure 3 depicts the comparison of the dynamics between the gKdV and the rescaled KdV equations for the initial datum obtained from (4.7) with $c=0.4$ and $x_{0}=0$. We observe that the timescales required for the distance $\left\|\Delta(t)\right\|_{H_{x}^{s}(\mathbb{R})}$ to reach $\mathcal{O}(1)$ are of $\mathcal{O}(10^{3})$ ($t\sim 4000$ in this example). In the critical case $k=4$,  we detect that $\nu(4)\sim 0.41$  and the timescales for the distance to reach $\mathcal{O}(1)$  are $t\sim\mathcal{O}(10^{2})$.   Within these large timescales (in view of the degree of the nonlinearity as discussed above), the observed deviation is primarily due to weak radiative effects exhibited by the gKdV dynamics, while the soliton’s initial amplitude and shape persist with only minor differences from the analytical rescaled soliton profile.
@@
Figure 3. Left and middle: Snapshots comparing the profiles of the solution of the gKdV equation with power nonlinearity (1.3) for $k=3$ (solid blue) against the profiles of the soliton (4.7) of the rescaled integrable KdV (4.6) with $\nu=0.47$ (red dashed). Right: The top panel shows the contour plot of the evolution of the initial datum (4.4) with $c=0.4$ and $x_{0}=0$ under gKdV. The dashed black line is the caustic of the rescaled soliton (4.7) with $\nu=0.47$. The bottom panel presents the evolution of the distance measured in the spaces $\mathcal{X}=H_{x}^{s}$, $s=0,1,2$, and $L_{x}^{\infty}$.
@@
Inspecting the evolution of the distance norms in Figures 2 and 4 for both the $k=3$ and $k=4$ examples, we observe a distinct dynamical pattern within the time interval of integration. The distance first reaches a local maximum at the end of the initial linear growth regime. Subsequently, the distance decreases, attaining a minimum, after which the distance resumes an almost linear growth. This feature can be explained as follows. The power nonlinearity affects the KdV soliton initial datum by introducing “curvature” to its caustic, along which the solitary wave moves at a significantly lower speed than the analytical KdV soliton (as seen in Figure 2 for $k=3$, prior to the caustic rotation). This feature is more clearly visualized when comparing the contour plot and evolution of norms in Figure 4 for $k=4$, where the curvature of the gKdV caustic is clearly visible. For the case $k=3$, the curvature of the gKdV caustic is very small relative to the time of evolution, causing the caustics to appear as almost coinciding. After the rotation, the caustic of the rescaled KdV soliton (dashed black lines in the relevant contour plots) acts as a secant line traced from $t=0$, which approximates the slightly curved caustics of the gKdV solitary waves over large timescales. We relate the evolution of the norms to the relative speeds of the caustics:

### v2

@@ Abstract.
The distance between the solutions to the integrable Korteweg-de Vries (KdV) equation and a broad class of non-integrable generalized KdV (gKdV) equations is estimated in appropriate Sobolev spaces. This family of equations includes, as special cases, the standard gKdV equation with power nonlinearities as well as weakly nonlinear perturbations of the KdV equation. For initial data and nonlinearity parameters of arbitrary size, we establish distance estimates based on a crucial size estimate for local gKdV solutions that grows linearly with the norm of the initial data. Consequently, these estimates predict that the dynamics of the gKdV and KdV equations remain close over long time intervals for initial amplitudes approaching unity, while providing an explicit rate of deviation for larger amplitudes. These theoretical results are supported by numerical simulations of one-soliton and two-soliton initial conditions, which show excellent agreement with the theoretical predictions. Furthermore, it is demonstrated that in the case of power nonlinearities and large solitonic initial data, the deviation between the integrable and non-integrable dynamics can be drastically reduced by incorporating suitable rotation effects via a rescaled KdV equation. As a result, the integrable dynamics stemming from the rescaled KdV equation may persist within the gKdV family of equations over remarkably long timescales.
@@ 1. Introduction and main results
At this point, it is important to highlight a fundamental distinction between the present approach and the one of [DZ02], particularly regarding weak nonlinearities. The methodology in [DZ02] utilizes integrability theory by combining the inverse scattering transform for the analysis of the reflection coefficient in the defocusing NLS case with detailed estimates for the perturbed non-integrable defocusing NLS model. On the other hand, our approach for deriving distance estimates between the integrable and non-integrable models remains generic across the class of nonlinearities described by (1.2) for the (non-integrable) gKdV family of equations. Since our framework is independent of the size of the initial data and other parameters, it captures both proximity and deviation between the solutions of the compared models without explicitly relying on the integrability theory for the KdV equation. A potential synthesis of the methodology in [DZ02] with the methods developed herein could yield even more precise distance estimates, presenting a formidable and challenging task for future investigations.
@@ 2. Proof of the size estimate
In the case of $U_{0}\in\mathcal{S}(\mathbb{R})$, the results of [Kat83] (see also [Tem69, Sjö70, BS75]) imply the existence of a unique solution in $U\in C_{t}^{\infty}([0,T_{0}];\mathcal{S}(\mathbb{R}))$, where $T_{0}>0$ depends only on the norm of $U_{0}$. For that smooth solution, let $n\in\mathbb{N}\cup\{0\}$ with $0\leq n\leq s$ and apply $\partial_{x}^{n}$ to equation (1.1). Then, multiply the resulting equation by $\partial_{x}^{n}U$ and integrate in $x$ while recalling that $U$ is real-valued to obtain
@@
For $0\leq j\leq n-2$, by the Sobolev embedding theorem,
$\left\|\partial_{x}^{n-j}(F(U))\,\partial_{x}^{j+1}U\right\|_{L_{x}^{2}(\mathbb{R})}\leq\left\|\partial_{x}^{j+1}U\right\|_{L_{x}^{\infty}(\mathbb{R})}\left\|\partial_{x}^{n-j}(F(U))\right\|_{L_{x}^{2}(\mathbb{R})}\leq c\left\|U\right\|_{H_{x}^{j+2}(\mathbb{R})}\left\|\partial_{x}^{n-j}(F(U))\right\|_{L_{x}^{2}(\mathbb{R})}$ (2.12)
@@
$\left\|\partial_{x}^{n-j}(F(U))\,\partial_{x}^{j+1}U\right\|_{L_{x}^{2}(\mathbb{R})}\leq c\left\|U\right\|_{H_{x}^{s}(\mathbb{R})}\left\|\partial_{x}^{n-j}(F(U))\right\|_{L_{x}^{2}(\mathbb{R})}.$ (2.13)
@@
Next, we extend the result to the case of Sobolev solutions. To this end, by density, given $U_{0}\in H^{s}(\mathbb{R})$, we let $\{U_{0,n}\}_{n\in\mathbb{N}}\in\mathcal{S}(\mathbb{R})$ be such that $\left\|U_{0}-U_{0,n}\right\|_{H^{s}(\mathbb{R})}\to 0$ as $n\to\infty$. Then, thanks to the result of [Kat83], there is a unique solution $U_{n}\in C_{t}^{\infty}([0,T_{0,n}];\mathcal{S}(\mathbb{R}))$ corresponding to $U_{0,n}$, where $T_{0,n}>0$ depends only on the norm of $U_{0,n}$. Hence, by estimate (1.21), which was established above for smooth solutions,
@@
$\leq 2\left\|U_{0,n}\right\|_{H^{s}(\mathbb{R})}+K\left\|U_{0}-U_{0,n}\right\|_{H^{s}(\mathbb{R})},\quad t\in[0,T_{n}].$
@@ 3. Proof of the distance estimates
$\left\|\Delta(t)\right\|_{L_{x}^{2}(\mathbb{R})}\leq\left\|\Delta(0)\right\|_{L^{2}(\mathbb{R})}+\int_{0}^{t}\left\|N(\tau)\right\|_{L_{x}^{2}(\mathbb{R})}d\tau.$ (3.6)
@@
$\left\|\Delta(t)\right\|_{L_{x}^{2}(\mathbb{R})}\leq\left\|\Delta(0)\right\|_{L^{2}(\mathbb{R})}+\sup_{\tau\in[0,t]}\Big[\left\|u(\tau)\right\|_{H_{x}^{1}(\mathbb{R})}^{2}+c_{k}\Big(\left\|U(\tau)\right\|_{H_{x}^{1}(\mathbb{R})}^{k+1}+B_{k}\sum_{j=1}^{k-1}\left\|U(\tau)\right\|_{H_{x}^{1}(\mathbb{R})}^{j+1}\Big)\Big]\,t.$ (3.10)
@@
$\sup_{t\in[0,T]}\left\|\Delta(t)\right\|_{L_{x}^{2}(\mathbb{R})}\leq\left\|\Delta(0)\right\|_{L^{2}(\mathbb{R})}+2^{k+1}\Big[\left\|u_{0}\right\|^{2}_{H^{2}(\mathbb{R})}+c_{k}\Big(\left\|U_{0}\right\|_{H^{2}(\mathbb{R})}^{k+1}+B_{k}\sum_{j=1}^{k-1}\left\|U_{0}\right\|_{H^{2}(\mathbb{R})}^{j+1}\Big)\Big]\,T.$ (3.11)
@@
$\left\|\Delta(t)\right\|_{H_{x}^{s}(\mathbb{R})}\leq\sqrt{2}\left\|\Delta(0)\right\|_{H^{s}(\mathbb{R})}+\sqrt{2}\int_{0}^{t}\left\|N(\tau)\right\|_{H_{x}^{s}(\mathbb{R})}d\tau.$ (3.13)
@@
$\left\|\Delta(t)\right\|_{H_{x}^{s}(\mathbb{R})}\leq\sqrt{2}\left\|\Delta(0)\right\|_{H^{s}(\mathbb{R})}+\sqrt{2}\sup_{\tau\in[0,t]}c_{s,k}\Big(\left\|u(\tau)\right\|_{H_{x}^{s+1}(\mathbb{R})}^{2}+\left\|U(\tau)\right\|_{H_{x}^{s+1}(\mathbb{R})}^{k+1}+B_{k}\sum_{j=1}^{k-1}\left\|U(\tau)\right\|_{H_{x}^{s+1}(\mathbb{R})}^{j+1}\Big)\,t$ (3.17)
@@
$\sup_{t\in[0,T]}\left\|\Delta(t)\right\|_{H_{x}^{s}(\mathbb{R})}$ $\leq\sqrt{2}\left\|\Delta(0)\right\|_{H^{s}(\mathbb{R})}+2^{k+\frac{3}{2}}\Big(\left\|u_{0}\right\|_{H^{s+1}(\mathbb{R})}^{2}+\left\|U_{0}\right\|_{H^{s+1}(\mathbb{R})}^{k+1}+B_{k}\sum_{j=1}^{k-1}\left\|U_{0}\right\|_{H^{s+1}(\mathbb{R})}^{j+1}\Big)\,T.$ (3.18)
@@ 4. Numerical simulations
(a)               (b)
Refer to caption Refer to caption
(c)               (d)
Refer to caption Refer to caption
Figure 1. The solution $U(x,t)$ of the gKdV equation (1.1) with monomial nonlinearity according to (4.1). (a) The supercritical case $k=5$ supplemented with KdV one-soliton initial data $u_{\text{s}}(x)$ with $c=0.01$ and $x_{0}=0$. (b) The critical case $k=4$ supplemented with mKdV one-soliton initial data $U_{\text{s}}(x)$ with $c=0.004$ and $x_{0}=0$. (c) Time evolution of the norms $\left\|\Delta(t)\right\|_{\mathcal{X}}$ of the difference between the above gKdV solutions and the corresponding KdV and mKdV one-solitons for $\mathcal{X}=H_{x}^{1}$, $L_{x}^{2}$ and $L_{x}^{\infty}$, for $k=5$. (d) The corresponding time evolution of the norms for $k=4$.
@@ 4.1. The gKdV with power nonlinearity
As a first example, we consider the initial datum $u_{\text{s}}(x)$ in (4.4) with $c=0.01$ so that, according to (4.5), $\left\|u_{\text{s}}\right\|_{H^{2}}\approx 0.1551$. We then evolve this datum under the supercritical gKdV equation (1.3) with $k=5$. The dynamics for this case are depicted in Figure 1, where panel (a) illustrates the contour plot of the solution $U(x,t)$ of the gKdV equation and panel (c) illustrates the time evolution of the norms $\left\|\Delta(t)\right\|_{\mathcal{X}}$. Evolving this datum over the time interval $t\in[0,\left\|u_{\text{s}}\right\|_{H^{2}}^{-1}]\simeq[0,6.5]$ and computing numerically the difference norm $\left\|\Delta(t)\right\|_{\mathcal{X}}$ for $\mathcal{X}=H_{x}^{s}(\mathbb{R})$, $s=0,1$, and $L_{x}^{\infty}(\mathbb{R})$, we observe the following: By Corollary 1, we have the theoretical bound $\left\|\Delta(t)\right\|_{\mathcal{X}}\lesssim\epsilon^{2}\approx 0.024$. The numerical simulations fulfill this proximity estimate as shown in panel (c) of Figure 1 and, in fact, for $t\in[0,80]$ yield $\left\|\Delta(t)\right\|_{L_{x}^{\infty}}\leq 0.0024$, which is an order of magnitude smaller than the theoretical bound. Moreover, the simulations exhibit a linear in time growth of the proximity error, in agreement with the theoretical predictions.
As a second example, we consider the second initial profile $U_{\text{s}}(x)$ in (4.4) with $c=0.004$ so that, by (4.5), $\left\|U_{\text{s}}\right\|_{H^{2}}\approx 0.8718$. We evolve the mKdV one-soliton data under the critical gKdV equation (1.3) with $k=4$. The corresponding dynamics are depicted in panels (b) and (d) of Figure 1; the format of the presentation is the same as in panels (a) and (c), respectively. For $t\in[0,\left\|U_{\text{s}}\right\|_{H^{2}}^{-2}]\approx[0,1.32]$, we compute numerically $\left\|\Delta(t)\right\|_{\mathcal{X}}$. By Corollary 1, we have $\left\|\Delta(t)\right\|_{\mathcal{X}}\lesssim\epsilon^{2}\approx 0.76$. Once again, the numerical results fulfill the theoretical estimates. Specifically, we find $\left\|\Delta(t)\right\|_{L_{x}^{\infty}}\leq 0.01$ for $t\in[0,100]$, which is again about an order of magnitude smaller than the theoretical bound. As in the first example, the growth of the distances in panel (d) of Figure 1 is in agreement with the theoretically predicted linear temporal growth of the proximity estimates.
@@
(a)               (b)               (c)
Refer to caption Refer to caption Refer to caption
(d)               (e)               (f)
Refer to caption Refer to caption Refer to caption
Figure 2. (a), (b), (d) and (e) Snapshots comparing the profile of the solution of gKdV with power nonlinearity (1.3) for $k=3$ (solid blue) against the profile of the solution of the integrable KdV (1.4) (red dashed). (c) Contour plot of the evolution of the initial datum $u_{\text{s}}(x)$ in (4.4) with $c=0.4$ and $x_{0}=0$ under gKdV. The dashed black line is the caustic of the analytical KdV soliton. (f) The evolution of the distance norms in $\mathcal{X}=H_{x}^{s}$, $s=0,1,2$, and $L_{x}^{\infty}$.
Panels (a), (b), (d) and (e) in Figure 2 illustrate snapshots at specific times comparing the profiles of the solutions. The gKdV solution is shown with the solid blue curve and the KdV soliton with the dashed red curve. Panel (c) shows the contour plot of the gKdV solution, together with the caustic of the corresponding analytical KdV soliton, depicted by the dashed black line. Panel (f) shows the evolution of the distance norm $\left\|\Delta(t)\right\|_{\mathcal{X}}$ with $\mathcal{X}=H_{x}^{s}(\mathbb{R})$, $s=0,1,2$, and $L_{x}^{\infty}(\mathbb{R})$.
According to Corollary 1, we expect proximity of the solutions only for very small timescales predicted by $\widetilde{T}_{s}=\mathcal{O}(\left\|u_{\text{s}}\right\|_{H^{2}}^{-3})$. This is already seen in the first snapshot at $t=1$ of panel (a), where we observe the beginning of the deviation of the numerical gKdV soliton from the analytical KdV one-soliton, manifested by the small increase in its amplitude and the onset of its radiation. In the snapshot at $t=10$ of panel (b), the radiation of the gKdV soliton is clearly observable and the KdV soliton is faster. These snapshots, as well as the one at $t=25$ of panel (d), depict profiles in the regime of the linear growth of the distances between the solutions. In particular, the snapshot at $t=25$ refers to a time close to the end of the linear growth regime of $\left\|\Delta(t)\right\|_{\mathcal{X}}$ shown in panel (f), manifested by the complete separation of the soliton profiles, as observed in the snapshot at $t=200$ of panel (e); the beginning of this separation at $t\sim 25$ is also visible in the contour plots of the solutions shown in panel (c).
@@
We detect the following dynamical feature: for a given nonlinearity exponent $k\geq 2$, there exists a scaling parameter $0<\nu(k)<1$ such that the soliton dynamics in the gKdV equation remain remarkably proximal, for significantly large time intervals, to the analytical rescaled soliton (4.7).   For gKdV with $k=3$, we detect a scaling parameter $\nu(3)\sim 0.47$. Figure 3 depicts the comparison of the dynamics between the gKdV and the rescaled KdV equations for the initial datum obtained from (4.7) with $c=0.4$ and $x_{0}=0$; the format of the presentation is the same as in Figure 2. We observe in panel (f) of Figure 3 that the timescales required for the distance $\left\|\Delta(t)\right\|_{H_{x}^{s}(\mathbb{R})}$ to reach $\mathcal{O}(1)$ are of $\mathcal{O}(10^{3})$ ($t\sim 4000$ in this example). In the critical case $k=4$, for which the relevant dynamics are shown in Figure 4, we detect that $\nu(4)\sim 0.41$  and the timescales for the distance to reach $\mathcal{O}(1)$  are $t\sim\mathcal{O}(10^{2})$.   Within these large timescales (in view of the degree of the nonlinearity as discussed above), the observed deviation is primarily due to weak radiative effects exhibited by the gKdV dynamics, while the soliton’s initial amplitude and shape persist with only minor differences from the analytical rescaled soliton profile.
(a)               (b)               (c)
Refer to caption Refer to caption Refer to caption
(d)               (e)               (f)
Refer to caption Refer to caption Refer to caption
@@
Figure 3. (a), (b), (d) and (e) Snapshots comparing the profiles of the solution of the gKdV equation with power nonlinearity (1.3) for $k=3$ (solid blue) against the profiles of the soliton (4.7) of the rescaled integrable KdV (4.6) with $\nu=0.47$ (red dashed). (c) Contour plot of the evolution of the initial datum (4.4) with $c=0.4$ and $x_{0}=0$ under gKdV. The dashed black line is the caustic of the rescaled soliton (4.7) with $\nu=0.47$. (f) The evolution of the distance measured in the spaces $\mathcal{X}=H_{x}^{s}$, $s=0,1,2$, and $L_{x}^{\infty}$.
@@
Inspecting the evolution of the distance norms in Figures 2 and 4 for both the $k=3$ and $k=4$ examples, we observe a distinct dynamical pattern within the time interval of integration. The distance first reaches a local maximum at the end of the initial linear growth regime. Subsequently, the distance decreases, attaining a minimum, after which the distance resumes an almost linear growth. This feature can be explained as follows. The power nonlinearity affects the KdV soliton initial datum by introducing “curvature” to its caustic, along which the solitary wave moves at a significantly lower speed than the analytical KdV soliton (as seen in panel (c) of Figure 2 for $k=3$, prior to the caustic rotation). This feature is more clearly visualized when comparing the contour plot and evolution of norms as seen in panels (c) and (f) of Figure 4 for $k=4$, where the curvature of the gKdV caustic is clearly visible. For the case $k=3$, the curvature of the gKdV caustic is very small relative to the time of evolution, causing the caustics to appear as almost coinciding. After the rotation, the caustic of the rescaled KdV soliton (dashed black lines in the relevant contour plots) acts as a secant line traced from $t=0$, which approximates the slightly curved caustics of the gKdV solitary waves over large timescales. We relate the evolution of the norms to the relative speeds of the caustics: